### v1

# VidMuse: A Simple Video-to-Music Generation Framework with Long-Short-Term Modeling

Abstract
In this work, we systematically study music generation conditioned solely on the video. First, we present a large-scale dataset comprising 190K video-music pairs, including various genres such as movie trailers, advertisements, and documentaries. Furthermore, we propose VidMuse, a simple framework for generating music aligned with video inputs. VidMuse stands out by producing high-fidelity music that is both acoustically and semantically aligned with the video. By incorporating local and global visual cues, VidMuse enables the creation of musically coherent audio tracks that consistently match the video content through Long-Short-Term modeling. Through extensive experiments, VidMuse outperforms existing models in terms of audio quality, diversity, and audio-visual alignment. The code and datasets will be available at https://github.com/ZeyueT/VidMuse/.
1 Introduction
Music is an essential element of video production, as it can enhance emotional feelings and convey the theme of the video content. Along with the development of social media platforms i.e., YouTube and TikTok, some studies [53, 12, 56] have shown that a piece of melodious music can vastly attract the audience’s attention and interest in watching the video. It thus leads to a great demand for video-to-music generation, which has drawn considerable attention from some researchers [16, 37, 70, 23, 33, 30].
Nevertheless, music creation for video is a challenging task, which requires both musical skills and deep understanding of the video semantics. It would be very time-consuming to produce a piece of suitable music for video in a hand-crafted manner. Therefore, it is desirable when we can automatically generate high-quality music for different genres of videos.
Currently, most of works [11, 34, 86, 22, 35, 35, 66] have made significant achievements, especially in text-to-music generation, but the video-to-music generation still remains to be further studied. Specifically, existing works on video-conditioned music generation mainly focus on specific scenarios, such as dance videos [44, 95], or on the symbolic music, i.e., MIDI [80, 16, 96, 37]. However, these works are unable to generate more diverse musical styles and are also difficult to generalize to various video genres. Moreover, Hong et al. [33] build a music–video retrieval dataset from YouTube-8M [3], albeit with limited video genres. Despite that there are also some prominent works [36, 70] employing multi-modal inputs to generate music for video, it is still worth studying whether it is possible to generate high-quality and visually consistent music solely based on the video input, while also generalizing to diverse genres of videos.
Motivated by this, we first construct a large-scale dataset termed V2M, equipped with a comprehensive benchmark to evaluate the state-of-the-art works thoroughly. The video-music pairs are collected from YouTube with various genres, e.g., movie trailers, advertisements, documentaries, vlogs, etc. In order to ensure the quality of our dataset, we establish a multi-step pipeline illustrated in Fig. 2 to systematically clean and pre-process data. The videos with low quality or composed of static images are filtered out. The proposed dataset contains three subsets: V2M-190K for pretraining, V2M-20K for finetuning, and V2M-bench for evaluation. We believe that V2M is able to facilitate the development of video-to-music generation.
Furthermore, on top of V2M, we propose a simple yet effective method, termed as VidMuse, to generate music only conditioned on the visual input. Instead of predicting the intermediate musical symbols such as MIDI or retrieving the music from the database, the proposed VidMuse, integrates both local and global visual clues to generate background music consistent with the video in an end-to-end manner. The core techniques in our method are a Long-Short-Term Visual Module and a Music Token Decoder. Specifically, the long-short-term visual module aims to learn the spatial-temporal representation of videos, which is the key to generating music aligned with the video. On the one hand, the long-term module models the entire video, capturing global context to understand the whole video. It contributes to the coherence of generated music at the video level. On the other hand, the short-term module focuses on learning the fine-grained clues at the clip level, which plays an important role in generating diverse music. The integration of two modules can improve the quality and visual consistency of generated music. In addition, the Transformer-based music token decoder is an autoregressive model, converting video embeddings obtained by Long-Short-Term Visual Module into music tokens. We formulate music generation as a task of next token prediction, which has been widely validated by the NLP community. The predicted music tokens are further decoded into the music signals by a high-fidelity neural audio compression model [20].
In summary, the main contributions of our work are as follows:
(a) We construct a large-scale video-to-music dataset, i.e., V2M, which contains about 190k video-music pairs with high quality, covering various genres and themes. To the best of our knowledge, this is the largest and most diverse dataset for this task, which can facilitate future research.
(b) We propose a simple yet effective method, VidMuse, for video-to-music generation. The proposed method integrates both local and global cues in the video, enabling the generation of high-fidelity music tracks that are not only musically coherent but also semantically aligned with the video content.
(c) We benchmark several state-of-the-art works against our method on V2M-bench with a series of subjective and objective metrics for thorough evaluation. As demonstrated in experiments, VidMuse achieve state-of-the-art performance on V2M-bench, outperforming existing models in terms of audio quality, diversity, and audio-visual consistency.
2 Related Work
In this section, we review the existing works related to video-to-music generation, which mainly fall into four categories:
Video Representation. Various methods have been proposed to learn the spatio-temporal representation  [48, 50, 75, 21, 6, 73, 91, 49] for videos. They aim to capture the contextual features of video frames, which is beneficial for video understanding. Recent advances primarily concentrate on developing video transformers, including ViViT [6, 73, 49]. These transformer-based methods achieve superior generalized performance on various video understanding tasks, such as video classification and temporal action localization. Among them, Tong et al. [73] extend masked autoencoders [29] from the image to the video, exhibiting the strong generalized performance in downstream tasks. Benefiting from the advance in multi-modal large language models, lots of works [91, 45, 59] of interactive video understanding have been proposed, which built upon the large language models (LLMs) [74, 92, 62, 71] and showcase the visual reasoning capabilities for video understanding.
Audio-Visual Alignment. Audio-visual alignment [5, 65, 68, 10, 26, 94, 87, 78, 82, 84] aims to align the feature between audio, vision in the semantics level, which plays a vital role in tasks of audio-visual understanding and generation. For example, CAV-MAE [26] is an audio-visual MAE by integrates the contrastive learning and masked modeling method. Currently, many works go beyond exploring audio-visual alignment. ImageBind [25] extends CLIP [61] to more modalities, including audio, depth, thermal, and IMU data, which paves the way for cross-modal retrieval and generation. In addition, Wu et al. [82] employ LLMs with multi-modal adaptors to support any modal data as input and output, showing strong capabilities in universal multi-modal understanding. These methods transcend audio-visual alignment, and dramatically advance the development of multi-modal representation learning.
Conditional Music Generation. Despite that there are lots of methods [18, 28, 34, 47, 57, 52, 54] studying unconditional music generation, in this paper, we mainly focus on reviewing the methods of conditional music generation, which are more related to our work. Many researchers [66, 4, 86, 22, 35, 11, 88, 14] make their endeavours on text-to-music generation. Similar to Stable Diffusion [63], these works [86, 22, 35, 66] try to adapt diffusion models for music generation. M2UGen [36] is a multi-modal music understanding and generation system that leverages large language models to process video, audio, and text. Video2Music[37] can generate music that matches the content and emotion of a given video. Moreover, the proposed V2Meow [70] not only can generate music that supports style control via text prompts. In contrast to previous video-to-music works [23, 37, 96, 70], our VidMuse utilizes a short-term module and a long-term module to model local and global visual clues in videos. As a result, it can generate high-fidelity music aligned with the video.
Video-to-music Datasets. Many multi-modal datasets [9, 58, 72, 42, 24, 93, 55, 69, 46, 3, 67, 31, 24] have been released, but there is still a lack of datasets for video-to-music generation. Hong et al. [33] construct the HIMV-200K with video-music pairs and aim to retrieve music for the video from the database. However, this dataset exhibits limited video genres and also suffers from the issue of data quality, as stated in  [96]. We observe that several works [80, 16, 96] aim to facilitate the MIDI music generation. However, we argue this musical form imposes restrictions on diversity for the music generation. Other datasets [95, 44] focus on generating music for dance videos only and have limited data size, which limits their applicability for general video-to-music models. As a result, this work constructs a large-scale video-to-music dataset where the music directly in wav format is diverse. We establish a rigorous pipeline for data collection and cleaning, which ensures the quality and diversity of our dataset. We expect the model can learn the music with more diverse forms from this dataset.
3 Dataset
In this section, we build a multi-step pipeline to clean and process source videos from YouTube to ensure data quality. After that, we construct a large-scale video-to-music generation dataset, i.e., V2M, with a benchmark. The constructed dataset stands out for its significant size, high quality and rich diversity, including a wide range of genres such as movie trailers, advertisements, documentaries, vlogs, etc. This comprehensive and diverse dataset aims to facilitate the video-to-music generation.
Figure 1: Dataset Construction. To ensure data quality, V2M goes through rule-based coarse filtering and content-based fine-grained filtering. Music source separation is applied to remove speech and singing signals in the audio. After processing, human experts curate the benchmark subset, while the remaining data is used as the pretraining dataset. The pretrain data is then refined using Audio-Visual Alignment Ranking to select the finetuning dataset.
3.1 Dataset Collection
To quickly collect a large scale of video-music pairs, we curate a series of query sets to retrieve corresponding videos from YouTube. In addition, we found that the music in the movie trailer usually showcases rich diversity and high quality. Therefore, we also aggregate a vast array of video information from the IMDb Non-Commercial Datasets [1], including video types, names, release dates, etc. Queries are formulated by appending “official trailer” to the titles of these selections and retaining the videos released after 2000, as videos from earlier periods are less likely to be of good quality. In the process of video crawling, we only keep the top 2 search results, resulting in a collection of around 400K videos, ranging from movies to documentaries.
3.2 Dataset Construction
The 400K raw videos may include many low-quality samples. To address this, we develop a series of rigorous steps to filter the undesirable data and obtain a clean set. We further process the source videos from the clean set, separating the irrelevant vocal tracks while preserving the music tracks. After that, the final kept videos with high quality can be directly used for training models. The overall pipeline of data processing is shown in Fig. 1.
Coarse Filtering. We design a rule-based filtering strategy for initial data screening. First, we perform illegal video and audio filters, which filter out the video without an audio track or a video track. Next, we apply a duration filter to filter out videos based on their duration, excluding those that are either too long (over 480 seconds) or too short (under 30 seconds). Finally, we we use a domain filter to examine metadata and exclude specific categories such as Interview, News, Gaming, as the background music in these genres typically lacks semantic alignment with the visual content.
Fine-grained Filtering. To further ensure the quality of our data, we conducted additional audio and visual analyses. For the audio analysis, raw videos may contain audio segments without music, such as speech, silence, etc. To ensure the final dataset consists of high-quality video-music pairs, we retain only those videos with a larger portion of music content. We utilize the sound event detection model PANNs [39], which provides frame-level event labels across the entire video to identify music events. Based on the observation from a subset of videos, we define two thresholds, i.e., a confidence threshold and a duration threshold for analyzing the music event. The confidence threshold is set at $0.5$, indicating an audio frame is considered a music event if the PANNs model predicts the probability of the “Music” label to be over $0.5$. The duration threshold of a music event requires that at least $50\%$ of the audio’s frames are classified as music events for the video to be considered valid.
For the visual analysis, some videos only consisting of static images will be removed. Specifically, we uniformly sample multiple temporal windows without overlap from the video. Within each window, we use Structural Similarity Index Measure (SSIM) [79] between the first frame and the last frame. By aggregating average SSIM values from all temporal windows, we remove the videos with average SSIM values lower than a threshold of 0.8, empirically.
Music Source Separation. Since the irrelevant human speech in videos poses a negative impact on music generation, we apply music source separation to process the videos. We employ Demucs [13] as the music source separation model to filter out the speech signals.
Audio-Video Alignment Ranking. ImageBind-AV [25] scores usually reflect the semantic correlation between the vision and audio modality. To construct a high-quality subset with better alignment, we compute the ImageBind-AV scores for all the data and rank them accordingly.
After filtering and ranking, we split the final videos into the training set, V2M, from all the paired data. The top 20K pairs are selected to form the finetuning subset, V2M-20K. In addition, we randomly sample 1,000 videos excluded from the training set. These 1,000 videos are then further evaluated by five human experts based on audio quality and the degree of audio-visual alignment. Ultimately, the top 300 high-quality videos are selected as a test set, termed as V2M-bench.
3.3 Data Analysis
The above data pipeline yields three data splits. Specifically, the training set comprises $\sim$190 video-music pairs, totaling 6,403 hours. The finetuning dataset consists of $\sim$20K pairs, totaling 597 hours. The benchmark dataset contains 300 pairs, with a cumulative duration of 9 hours. Fig. 1(a) showcases the category and genre distribution of our training data, highlighting its comprehensive diversity. This diversity ensures a rich and varied dataset for the model training. As shown in Fig. 1(b), we compare with other related datasets, demonstrating its advantage in data scale. Through Fig. 1(c) and Fig. 1(d), we observe that the duration distribution across the training set, finetuning set, and test set is highly uniform, with the duration of most videos concentrated around 100 seconds.
Dataset Necessity. Some existing video-music pair datasets have been released [33, 80, 16, 96, 95, 44], buts some of them [80, 16, 96] aim to facilitate MIDI music generation, which limits the form of music. Datasets like [95, 44] focus on generating music for dance videos only and have limited data size. The dataset constructed by [33] includes video-music pairs but exhibits limited video genres and suffers from data quality issues. Motivated by this, we develop a multi-step pipeline and curate a large-scale dataset V2M for the video-to-music generation.
3.4 Evaluation Metrics
To quantitatively evaluate the effectiveness of our model, we employ a series of metrics to assess different models in terms of quality, fidelity, and diversity for generated music. Besides, we also examine the alignment degree between the video and generated music. We believe these metrics can thoroughly evaluate the performance of different methods.
Frechet Audio Distance (FAD) [^1] https://github.com/haoheliu/audioldm_eval [38] is a reference-free evaluation metric for assessing audio quality. Similar to Frechet Image Distance (FID)[32], it compares the embedding statistics of the generated audio clip with groundtruth audio clips. A shorter distance usually denotes better human-perceived acoustic-level audio quality. However, this metric cannot reflect semantic-level information in audio. We report the FAD based on the VGGish[31] feature extractor.
Frechet Distance (FD).33footnotemark: 3 measures the similarity between generated samples and target samples in audio generation fields. It’s similar to FAD but uses a PANNs feature extractor instead. PANNs[39] have been pre-trained on AudioSet[24], one of the largest audio understanding datasets, thus resulting in a more robust metric than FAD.
Kullback-Leibler Divergence (KL).33footnotemark: 3 reflects the acoustic similarity between the generated and reference samples to a certain extent. It is computed over PANNs’ multi-label class predictions.
Density and Coverage [^2] https://github.com/clovaai/generative-evaluation-prdc [60] measures the fidelity and diversity aspects of the generated samples. Fidelity measures how closely the generated samples match the real ones, while diversity assesses whether the generated samples capture the full range of variation found in real samples. We use CLAP[83] embeddings for manifold estimation.
Imagebind Score [^3] https://github.com/facebookresearch/ImageBind [25] assesses to what extent the generated music aligns with the videos. Despite the fact that Imagebind extends the CLIP to six modalities, we only use the branches of audio and vision. Since we use ImageBind to filter out video-audio pairs with a low matching score during dataset construction, the ImageBind score is naturally used in our evaluation. We acknowledge that imageBind is not specifically trained on music data, but there is no better choice currently.
4 Method
Figure 2: Overview of the VidMuse Framework.
In this section, we elaborate on the details of our proposed VidMuse. Our approach leverages Long-Short-Term modeling to generate music that is aligned with video content. Specifically, the Short-Term Module captures localized visual features to ensure the generated music aligns with local variations in the video, while the Long-Term Module models the entire video, providing global context to guide the Short-Term Module in generating more visually consistent music. The experiments in Sec. 5 demonstrate the effectiveness of our proposed method.
4.1 Overview
We illustrate the VidMuse framework’s pipeline in Fig. 2, which converts visual content to music using a structured four-module architecture. The components include (1) a Visual Encoder, (2) a Long-Short-Term Visual Module, (3) a Music Token Decoder, and (4) an Audio Codec Decoder.
Given an input video X, the Visual Encoder extracts feature representations $\bar{\textbf{X}}$. These features are then fed into the Long-Short-Term Visual Module. Specifically, the Short-Term Module captures localized details, while the Long-Term Module models the entire video, providing global context, detailed in Sec. 4.3. This results in the feature representation Z through multi-head cross-attention. The feature Z is then processed by the Music Token Decoder to generate music tokens $\bar{\textbf{Y}}$. Finally, these music tokens are decoded by the Audio Codec to produce the final music $\bar{\textbf{A}}$.
4.2 Visual Encoder
The Visual Encoder processes input video frames, denoted by $\textbf{X}\in\mathbb{R}^{N\times C\times H\times W}$, where $N$ represents the frame count and each frame has dimensions $C\times H\times W$. This component extracts feature representations $\bar{\textbf{X}}\in\mathbb{R}^{N\times P\times D}$ from the video frames. Here, $P$ and $D$ denote the sequence length, including the class token, and the dimensionality of feature vectors, respectively [19, 61, 73, 6].
4.3 Long-Short-Term Visual Module
To capture both local and global visual clues, we model the visual features through the Long-Short-Term Visual Module. Specifically, we evenly samples $N_{l}$ frames from $\bar{\textbf{X}}$ to obtain global feature representations $\bar{\textbf{X}}_{l}\in\mathbb{R}^{N_{l}\times P\times D}$, and extract a continuous segment of $N_{s}$ frames from $\bar{\textbf{X}}$ to form local feature representations $\bar{\textbf{X}}_{s}\in\mathbb{R}^{N_{s}\times P\times D}$. The representations are then refined through Long-term Modeling and Short-term Modeling, respectively, resulting in long-term features $\bar{\textbf{F}}_{l}\in\mathbb{R}^{N_{l}\cdot\text{$N_{H}$}\times D}$ and short-term features $\bar{\textbf{F}}_{s}\in\mathbb{R}^{N_{s}\cdot\text{$N_{H}$}\times D}$, where $N_{H}$ represents the number of attention heads.
LST-Fusion integrates long-term and short-term features by utilizing the cross-attention mechanism in the form of Query ($\mathbf{q}$), Key ($\mathbf{k}$), and Value ($\mathbf{v}$), which is formulated as:
$\mathbf{q}={\bar{\textbf{F}}_{s}}\mathbf{W}_{q},\ \ \ \mathbf{k}=\bar{\textbf{F}}_{l}\mathbf{W}_{k},\ \ \ \mathbf{v}=\bar{\textbf{F}}_{l}\mathbf{W}_{v},$ (1)
$\textbf{Z}^{\prime}=\text{softmax}(\mathbf{q}\mathbf{k}^{T}/\sqrt{C/h})\mathbf{v}+\bar{\textbf{F}}_{s},$ (2)
where $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ $\in\mathbb{R}^{C\times(C/h)}$ denote trainable parameters, with $C$ representing the embedding dimension and $h$ the number of heads. This mechanism allows the model to query global information rather than generating music based solely on local visual features. It guarantees that the generated music is more consistent with the overall video content.
After the cross-attention, a linear projection layer reconfigures $\textbf{Z}^{\prime}$ to align with the Music Token Decoder’s requirements:
$\textbf{Z}=Linear(\textbf{Z}^{\prime}).$ (3)
The $Linear$ that is a fully connected layer projects $\textbf{Z}^{\prime}$ into a space characterized by $\textbf{Z}\in\mathbb{R}^{N_{s}\cdot N_{H}\times M}$, where $M$ corresponds to the intermediate layer dimension within the Music Token Decoder’s transformer architecture.
The Long-Short-Term Visual Module integrates short-term and long-term features of the video and yields a representation Z, ready for the generation of music tokens. The experiment results in Sec. 5 show the efficacy of this module.
4.4 Music Token Decoder
Music and language are both sequential and structured. We thus can try to adapt LLMs to music [76, 34, 15, 4, 11]. To apply LLM techniques to the audio field, several studies have utilized different audio codecs to represent audio information as discrete tokens [90, 40]. These codecs allow audio segments to be encoded into a matrix with dimensions $[K,T_{s}]$, where $K$ is the number of codebooks and $T_{s}$ is the number of sequence timesteps. Some researchers have also investigated various ways to arrange these codebooks [90, 77, 11, 27, 85, 41].
From the previously described steps, we obtain the feature set $\textbf{Z}\in\mathbb{R}^{N_{s}\cdot N_{H}\times M}$, which is then input into the Music Token Decoder’s transformer $\mathcal{G}$ with a linear classification head. This transformation stage produces an output tensor $\bar{\textbf{Y}}\in\mathbb{R}^{K\times S}$ as music tokens, where $K$ is the number of codebooks and $S$ denotes the sequence steps in the musical piece:
$\bar{\textbf{Y}}=Linear(\mathcal{G}(\textbf{Z})),$ (4)
where $\mathcal{G}$ aims to convert video features into a sequence of music features.
4.5 Optimization
Given the codec $\mathcal{C}$, we denote the Audio Encoder as $\mathcal{C}_{\text{encode}}(\cdot)$ and Audio Decoder as $\mathcal{C}_{\text{decode}}(\cdot)$, respectively. The process of encoding the ground truth audio A into its tokenized representation Y can be represented as
$\textbf{Y}=\mathcal{C}_{\text{encode}}(\textbf{A}).$ (5)
The audio decoder $\mathcal{C}_{\text{decode}}(\cdot)$ is utilized solely during the inference stage to decode music tokens into music. In the training stage, an input video X passes through a sequence of operations to generate a predicted representation $\bar{\textbf{Y}}$. With the target representation, Y, is derived by encoding the ground-truth audio A by eq. 5. The objective of our model is to minimize the divergence between $\bar{\textbf{Y}}$ and Y, formalized as the loss function:
$\mathcal{L}=\mathcal{L}(\textbf{Y},\bar{\textbf{Y}}),$ (6)
where $\mathcal{L}$ is cross-entropy loss in practice.
4.6 Inference
When predicting music on videos of arbitrary length, maintaining music consistency and coherence is particularly important. However, it leads to a significant challenge on computational resources due to the quadratic dependency of transformers-based models on sequence length [89, 7]. To address this problem, similar to methods used in previous works [81, 7], we adopt a sliding window approach for inferring the whole video.
During inference, given an input video X with the length of $L$, we set the sliding window length as $L_{s}$ and the length of the window’s overlap as $O$. The value of $t$ is initially set to 00. While $t+L_{s}\leq L$, the inference steps can be summarized as follows: 1) We use the visual encoder to extract feature embeddings $\bar{\textbf{X}}_{l}$ for a given video X. 2) Based on $\bar{\textbf{X}}_{l}$, we collect the feature embeddings within the time window of $[t:t+L_{s}]$ to obtain $\bar{\textbf{X}}_{s}$. 3) Conditioned on $\bar{\textbf{X}}_{l}$ and $\bar{\textbf{X}}_{s}$, the music token decoder predicts the music segment $\bar{\textbf{Y}}$ at the time window of $[t:t+L_{s}-O]$ after the long-short-term visual module. 4) The music tokens $\bar{\textbf{A}}$ are decoded from $\bar{\textbf{Y}}$ by a audio decoder: $\bar{\textbf{A}}=\mathcal{C}_{\text{decode}}(\bar{\textbf{Y}})$. 5) The slide the window moves forward, i.e., $t=t+L_{s}-O$, and we repeat steps 2 to 4 until the end of the video.
After finishing above steps, we can concatenate all musical segments to form a cohesive piece of music that aligns in duration with the video.
5 Experiments
In this section, we first elaborate on the implementation details of our experiment and then present the main results for both our method and the baseline methods using objective metrics and a subjective user study. Then, we demonstrate that our VidMuse exhibits impressive performance on all metrics. Furthermore, we conduct a comprehensive ablation study to explain our design choices and their impact on the performance.
5.1 Implementation details
We choose CLIP [61] as the visual encoder (refer to experiments in Appendix A). Since this work does not focus on audio encoding and decoding, we use Encodec [13] for 32 kHz monophonic audio as our default compression model and use the pretrained transformer model proposed in MusicGen [11]. The training stage utilizes the AdamW optimizer [51] for 22,000 steps with a batch size of 240 samples. The hyperparameters are set to $\beta_{1}=0.9$, $\beta_{2}=0.95$, with a weight decay of 0.1 and gradient clipping at 1.0. A cosine learning rate schedule is employed, incorporating a warm-up phase of 4,000 steps, and an exponential moving average decay of 0.99. The training is conducted on 48 H800 GPUs, with float16 mixed precision. This setup leverages $190$K samples for pre-training and $20$K samples for fine-tuning. A top-k strategy is applied for sampling, retaining the top $250$ tokens with a temperature setting of $1.0$. In the inference stage, we set the sliding window size as $30$s, and the length of the window’s overlap as $5$s.
5.2 Main Results
Table 1: Comparison with naive baselines and state-of-the-art methods.
     Methods               Metrics
[…]
  GT $0.000$ $0.000$ $0.000$ $1.167$ $1.000$ $0.241$
Caption2Music $1.081$ $40.199$ $2.485$ $0.378$ $0.486$ $0.191$
Video2Music [37] $1.782$ $144.881$ $18.722$ $0.103$ $0.023$ $0.136$
CMT [16] $1.220$ $85.704$ $8.637$ $0.080$ $0.070$ $0.124$
M2UGen [36] $0.997$ $52.246$ $5.104$ $0.608$ $0.433$ $0.181$
VM-NET [33] $0.899$ $67.480$ $6.252$ $0.986$ $0.383$ $0.147$
VidMuse 0.738 36.171 2.369 1.175 0.710 0.207
We benchmark several state-of-the-art methods, serving as baselines to compare with our method: 1) Caption2Music serving as a naive baseline, which employs the method [2] to extract the video captions and then outputs the music by feeding captions into MusicGen [11]. 2) Video2Music [37] and 3) CMT [16] which botch focus on predicting MIDI notes [64] from videos while our method directly generates music signals. 4) M2UGen [36], a strong baseline, which leverages a language model to connect vision and language, then use MusicGen [11] to generate music from language. 5) VM-NET [33], essentially different from the above methods, retrieves the piece of music from the database for a given video, while other methods predict music by learning from video-music pairs.
In Table 1, VidMuse integrating both global and local visual clues exhibits impressive performance on all metrics. Specifically, compared with Video2Music [37] and CMT [16], VidMuse shows the superiority in the diversity of generated music based on the metrics of density or coverage, which justifies the advantage of directly predicting music signals compared with MIDI notes. In addition, our method even outperforms the strong competitors, i.e., M2UGen. It proves that predicting music directly based on video input can also achieve better performance. Furthermore, compared with a retrieval-based method i.e., VM-NET [33], VidMuse achieves a higher Imagebind score, indicating that the music generated by the learning-based method is more consistent with the video semantics.
5.3 User Study
In our user study, we randomly sample 600 video-music pairs from the benchmark to conduct an A/B test, which is a widely used subjective evaluation method in the music field [17, 88], across five music generation methods: CMT, M2UGen, Caption2Music, Ground Truth (GT), and VidMuse. The test was distributed among 40 participants, ensuring each method was compared against another 60 times. The evaluation criteria are four-fold: 1) Audio quality: Refers to the sound quality of the audio track; 2) Video-music alignment: Assesses how well the music matches the visual content, e.g., a scene showing a woman crying should ideally be paired with music that sounds sad; 3) Musicality: Measures the aesthetic appeal of music, independent of audio quality, e.g., a piece of music has high audio quality but is out of tune has poor musicality; 4) Overall assessment: A comprehensive evaluation that considers all aspects of the music.
(a) Audio quality.
(b) Video-music alignment.
(a) Musicality.
(b) Overall assessment.
Figure 4: Extended A/B test user study results. Please refer to Sec. 5.3 where we introduce four criteria, i.e., audio quality, video-music alignment, musicality and overall assessment.
Participants are asked to choose the better sample for each criterion. The Fig. 4 illustrates the statistical results of the user study, where the value at matrix[$i$][$j$] ranges from 00 to $100$, indicating the % of times listeners preferred the method correspond to the $i$-row compared to the method correspond to the $j$-column. For example, in Fig. 4 (c), the value of matrix[$2$][$4$] represents that VidMuse outperforms CMT in $77\%$ of cases in terms of Musicality. Across all evaluation criteria, our method surpasses other methods in more than half of the comparisons, except when compared to the ground truth. Overall, these results robustly validate VidMuse’s effectiveness from the subjective evaluation.
5.4 Ablation Studies
In this section, we conduct comprehensive ablation studies to determine the optimal hyperparameters and design choices. First, we examine the impact of different model design choices on the generation quality. Second, we investigate the influence of the different video sampling rates and the duration of video segments in the Short-Term module on the results. After that, we explore the impact of the model size on performance. More ablation results are in the Appendix A.
Justification of Design Choices. To validate the impact of different model design choices on our generation results and verify the effectiveness of our method, we first design two modules: a short-term modeling module (STM) and a long-term modeling module (LTM). VidMuse-STM aims to ablate the contribution of STM by using only STM without LTM while VidMuse-LTM utilizes only LTM. We present the results in Tab. 2. Based on these results, we learn an insight that local information plays a more importance role in the generation. Second, we implement two variants with Cross-Attention with learnable Queries (CAQ) in our framework. Specifically, CAQ_SL first uses a CAQ where $K$ and $V$ are short-term features and then uses a CAQ where $K$ and $V$ are long-term features. CAQ_LS do it in an opposite order. As shown in Tab. 2, VidMuse outperforms two variants. It seems that the learnable query requires a deeper architecture, as demonstrated by DETR [8] and Q-Former [43], but this manner reduces the model’s efficiency. Furthermore, we also try a baseline using Slowfast [21] and find its performance degenerates in Tab. 2. This may be because the fast path uses a higher frame rate in inputs than the slow path, but their temporal receptive fields are identical. Thus, Slowfast lacks guidance from global clues.
Exploration on model inputs. To explore the impact of different video sampling rates and the duration of video segments in the Short-Term module on performance, we conducted ablation studies on input FPS and short-term segment duration, detailed in Table 3. To intuitively assess the effectiveness of different settings, we employ an Average Rank (AR) metric. The AR metric ranks the results for a metric across all methods within the same table. The ranking result is from $1$ to $N$ (equals to the number of methods within the table), where $1$ is the best and $N$ is the worst. We eventually obtain AR results by averaging the ranking results for all metrics. Note that the AR results cannot be compared across different tables, since this metric is designed to clearly showcase the dominance for each method within one table. From Table 3, we observe that increasing both FPS and duration tends to enhance model capabilities, suggesting that denser frame sampling yields a more detailed video representation, thereby improving music generation. Nevertheless, to balance computational costs and performance, we use a 30-second duration at 2 FPS as our optimal setting.
The impact of model size on performance. For the architecture of the music token decoder, we test three variants with different scales: small ($455$M), medium ($1.9$B), and large ($3.3$B). The medium-sized VidMuse-M is recognized as the default setting if not stated, as it achieves the best results across all metrics, outlined in Table 4. We observe the larger VidMuse-L model fails to deliver proportional improvements. This discrepancy can be partly attributed to limited GPU resources. The model’s performance is near saturation, and the larger model does not exhibit significant advantages. As a result, VidMuse-M is recognized as the optimal choice, achieving a trade-off between performance and efficiency.
Table 2: Ablation studies on design choices.
     Methods               Metrics
[…]
  VidMuse-STM $0.898$ $45.752$ $4.915$ $1.124$ $0.470$ $0.196$
VidMuse-LTM $0.858$ $53.907$ $16.074$ $1.439$ $0.406$ $0.205$
VidMuse-CAQ_SL $0.843$ $48.940$ $3.733$ $0.947$ $0.547$ $0.163$
VidMuse-CAQ_LS $0.919$ $45.335$ $2.917$ $0.562$ $0.720$ $0.181$
VidMuse-Slowfast $1.511$ $84.683$ $10.029$ $0.266$ $0.285$ $0.118$
VidMuse 0.738 36.171 2.369 1.175 0.710 0.207
Table 3: Ablation studies on Video Duration and FPS.
     Duration(s) FPS                 Metrics
[…]
  5 2 $0.820$ $51.101$ $4.117$ $1.430$ $0.74$ $0.148$ $7.00$
15 2 $0.849$ $41.131$ $2.709$ $1.406$ $0.803$ $0.181$ $5.33$
30 2 $0.843$ $41.354$ $\underline{2.413}$ $1.487$ $\underline{0.840}$ 0.193 $3.67$
5 4 0.800 $51.540$ $4.343$ $1.271$ $0.787$ $0.145$ $7.17$
15 4 $0.830$ $41.154$ $2.562$ $1.278$ $0.823$ $0.176$ $5.17$
30 4 $0.849$ $\underline{40.032}$ $2.418$ $\underline{1.538}$ 0.843 0.193 $\underline{2.84}$
5 8 $\underline{0.819}$ $50.667$ $4.069$ $1.515$ $0.743$ $0.153$ $5.67$
15 8 $0.857$ $42.106$ $2.790$ $1.476$ $0.753$ $\underline{0.187}$ $6.00$
30 8 $0.824$ 38.942 2.299 1.573 0.843 $0.180$ 2.17
Figure 5: Qualitative Comparison results on sound spectrograms produced by different methods.
Table 4: Ablation studies on model size.
     Models Params               Metrics
[…]
  VidMuse-S $455$M $0.854$ $57.317$ $4.695$ $1.349$ $0.433$ $0.165$
VidMuse-M $1.9$B 0.843 41.354 2.413 1.487 0.840 0.193
VidMuse-L $3.3$B $0.873$ $53.955$ $4.041$ $1.330$ $0.453$ $0.170$
5.5 Qualitative Analysis
Our qualitative analysis, presented in Fig. 5, underscores specific limitations in the methods employed by CMT and M2UGen. For CMT, the method’s reliance on extracting only low-level visual cues for music generation, particularly in creating symbolic music, leads to pronounced discontinuities. As illustrated in Fig. 5, “Abrupt gaps” occur, especially in darker video segments where the model fails to predict symbolic music notes, resulting in silence.
M2UGen’s use of LLMs and text embeddings to translate visual content into music affects its precision in capturing detailed visual nuances, leading to a noticeable lack of diversity, evidenced by repetitive musical themes. Additionally, M2UGen’s limitation to 30-second music generation segments requires the division of longer videos into clips for separate processing. This approach results in abrupt transitions between segments after stitching, which compromises the musical continuity.
The last row of Fig. 5 showcases our Long-Short-Term (LST) approach, highlighting its ability to produce globally diverse music that captures the essence of the ground truth. This method ensures our music generation remains contextually rich and aligned with the video.
6 Conclusion
In this work, we build a rigorous pipeline to collect high-quality and diverse video-music pairs, curating a comprehensive dataset V2M. Then, we propose VidMuse, a simple yet effective method for video-to-music generation. Our approach utilizes a Long-Short-Term approach to capture both local and global visual clues in the video, allowing for the generation of contextually rich and musically diverse outputs. To validate our method, we benchmark a series of state-of-the-art methods as baselines to compare with VidMuse. Through comprehensive quantitative studies and qualitative analyses, our method has demonstrated its superiority over the existing methods.
Limitations. Our work achieves a significant advancement in video-to-music generation, but it still has some limitations. First, the current implementation relies on a EnCodec model [20] trained on a dataset that restricts the system’s capacity to a 32kHz audio sampling rate. This codec exhibits a noticeable reconstruction loss in the audio signal, potentially lowering the quality of the generated music. Second, training large models requires substantial computational resources. Our future work aims to overcome these limitations by integrating advanced codec technologies to enhance audio reconstruction fidelity, and exploring a more efficient method design.
Appendix A Additional Experiments
Additional experiments focusing on pretraining effects, visual encoders, and codebook patterns are provided in the appendix. These parts provide insight into the decision-making process for selecting the experimental configurations within the VidMuse framework.
Pretraining Effect. Our ablation study on the effects of the data scale during finetuning, as detailed in Table A1, highlights a balance between data size and model performance. Despite not performing best in all the metrics, the model finetuned with 20k pair data emerges as our choice. The 20k data offers a compelling trade-off: it significantly improves performance across key metrics without requiring the extensive computational resources that larger datasets demand.
Visual Encoder. We conduct an ablation experiment to study the impact of visual encoders. As shown in table A2, we observe that VidMuse is robust in processing visual information for music generation. To trade-off between computational costs and quality of generation, we select CLIP [61] as the default visual encoder if not specified.
Codebook Pattern. The exploration of codebook interleaving patterns has attracted attention from researchers across several domains [90, 77, 11, 85, 41]. In our ablation study focusing on the patterns, we find that while the Parallel and Vall-E [77] patterns align with the findings for text-to-music generation in MusicGen [11], the flattened codebook pattern does not consistently exceed the performance of the delay pattern in tasks of generating music from video. The delay pattern, notable for its relatively low computational cost, is therefore selected for our implementation. The results of this study are presented in table A3.
Table A1: Ablation studies on the ratio of finetuning data.
     Finetuning Data             Metrics
[…]
  0 $0.750$ 35.788 $18.847$ $1.088$ $0.653$ $0.194$
10k $0.792$ $35.824$ $2.383$ $1.135$ $0.653$ $0.206$
20k 0.738 $36.171$ 2.369 $1.175$ $0.710$ 0.207
40k $0.771$ $36.521$ $2.476$ $1.150$ 0.716 $0.195$
60k $0.760$ $36.444$ $20.519$ 1.213 $0.706$ $0.197$
Table A2: Ablation studies on video encoders.
     Encoders               Metrics
[…]
  ViViT [6] $0.822$ $37.167$ $1.962$ $1.433$ $0.870$ $0.194$
VideoMAE [73] $0.778$ $37.900$ $19.388$ $1.074$ $0.663$ $0.209$
CLIP [61] $0.843$ $41.354$ $2.413$ $1.487$ $0.840$ $0.193$
ViT [19] $0.876$ $50.427$ $4.823$ $1.081$ $0.503$ $0.195$
(a) The distribution of video genres in our dataset.
(b) Comparisons with different datasets.
(c) Duration distributions of pretraining and finetuning data.
(d) Duration distribution of our benchmark data.
Figure A1: The statistics of our dataset.
Table A3: Ablation studies on codebook pattern.
     Patterns               Metrics
[…]
  Parallel $0.921$ $68.603$ $18.243$ $0.562$ $0.183$ $0.166$
Flatten 0.819 $52.931$ $4.260$ $1.351$ $0.500$ 0.201
Delay $0.843$ 41.354 2.413 1.487 0.840 $0.193$
Vall-E $0.866$ $57.286$ $4.681$ $1.148$ $0.354$ $0.189$